MATH 367, spg 98 PROJECT 3: Normal Approximation Due: Mon Apr 6

NAME:

Problem: Use the procedure used in Project 3 to compare the following distributions:

1. a Binomial (50,.2) against a normal with mean 50(.2) = 10 and standard deviation \sqrt{50(.2)(.8)} = 2.8284.

2. a Binomial (100,.2) against the appropriate Normal.

3. a Poisson with parameter 10 against the appropriate Normal.

4. a Poisson with parameter 100 against the appropriate Normal. For each item, plot both distributions together, look at the error (the absolute value of the difference) and tell me what is the maximum error.

Before you plot, figure out what is the relevant range and set the index counter accordingly.